OpenStudy (anonymous):

If p(2n-1,n) : p(2n+1, n-1) = 22:7 find "n"

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

so we have\[\huge \frac{\frac{(2n-1)!}{(n-1)!}}{\frac{(2n+1)!}{(n+2)!}}=\frac{22}{7} \]right?

OpenStudy (anonymous):

or\[\frac{n(n+1)(n+2)}{2n(2n+1)}=\frac{22}{7}\]